# Lost in Space: Optimizing Tokens for Grammar-Constrained Decoding

### v1

@@ Abstract
General-purpose language models are trained to produce varied natural language outputs, but for some tasks like annotation or classification we need more specific output formats. LLM systems increasingly support structured output, sampling tokens according to a grammar, which enforces a format but which can also reduce performance. We ask whether there are systematic differences between grammars that appear semantically similar to humans. To answer this question, we test four popular model families with five token formats on four NLP benchmarks. All models perform most accurately when instructed to classify with real numbers. Performance also improves by 5%-10% when models are instructed to return tokens incorporating leading whitespace, which we find can help models avoid structural deficiencies in subword token representations. Format-based differences are largest for smaller models that are often used for local laptop-scale inference. We present best practices for researchers using language models as zero-shot classifiers with structured output.
Lost in Space: Optimizing Tokens for Grammar-Constrained Decoding
@@ 1 Introduction
One of the most promising applications of LLMs is parsing natural language into structured data (Radford et al., 2018; Brown et al., 2020). LLMs are increasingly used to structure data by academics in fields as diverse as literary studies (Bamman et al., 2024), legal studies (Gray et al., 2023; Hamilton, 2023), finance (Holter and Ell, 2023), and computational biology (Hou and Ji, 2024). It is likewise prevalent outside of academia, with Veselovsky et al. (2023) estimating up to 46% of Mechanical Turk requests involve LLMs.
Prompt: Rate the similarity of these words from 1 to 5: train and underground. Llama 3.1 8b: Digits-Only Set Space-Digits Set 12345 Token _1 _2 _3 _4 _5 12345 Rendered 12345 ↓                                            ↓ 3                                            5 Llama 3.1 picks a different token when given choices incorporating whitespace.
@@
Figure 1: The Llama 3.1 provides different answers to a question from the MEN Test Collection when constrained to produce digits from 1-5 with and without a leading space. Both sets of inputs and outputs will look indistinguishable to humans. Generation was run with a temperature of 0.
@@
Users don’t want to waste queries on useless or difficult-to-interpret output. Prior research has proposed competing methods for guaranteeing valid output, such as reinforcement learning from human feedback (RLHF) and instruction tuning (Ouyang et al., 2022). These methods are time consuming and expensive to implement, and are impractical for the variable and rapid-turnaround scenarios typical of LLM usage.
@@
An increasingly popular method for ensuring output consistency without fine-tuning is selectively sampling tokens according to formal grammars (Shin et al., 2021; Wang et al., 2023). Structured output, or grammar-constrained decoding (GCD), has become a regular feature included with LLM server software (Gerganov, 2024; Microsoft, 2024; Rickard, 2024). GCD works by modifying LLM inference algorithms to selectively mask certain tokens according to some formal grammar. Some servers allow users to specify grammars with regular languages (Willard and Louf, 2023; Microsoft, 2024), while others allow users to write context-free grammars in Backus–Naur form (BNF), a notation for conveniently expressing rewrite rules (Chomsky, 1956; Knuth, 1964; Gerganov, 2024). But despite this popularity, using GCD appears to degrade task accuracy (Willard and Louf, 2023; Park et al., 2024; Beurer-Kellner et al., 2024; Geng et al., 2023; Le et al., 2024). We present an example of this degradation in Figure 1. Why is this the case, and how can we mitigate it?
@@
We investigate this phenomenon by comparing the performance achieved when using one of five distinct token sets across four different data labeling scenarios with four families of pre-trained language models (Dubey et al., 2024). We present three key contributions: first, we note all models perform best when guided to only emit numbers in the closed interval $[0,1]$. Second, we find most models benefit from returning tokens incorporating leading whitespace, a key finding building on recent research emphasizing the critical nature of tokens incorporating leading whitespace among other syntactic tokens (Pimentel and Meister, 2024; Oh and Schuler, 2024). We hypothesize ill-formed subword token representations are the basis of this preference for otherwise semantically identical tokens, and find supporting evidence in the embedding weights of our models of interest. We finally propose a set of strategies for selecting target tokens when specifying grammars for GCD libraries. These strategies can help achieve better performance for structured outputs from general-purpose models, and are particularly valuable for users of smaller, less computationally intensive models.
@@ 2 Related Work
Grammar-model misalignment is a recognized issue with grammar-constrained decoding algorithms. Existing approaches involve modifying the decoding system by adding additional constraints. Geng et al. (2023) propose penalizing the model for emitting short strings, a solution unfortunately unlikely to extend to context-free grammars when desiring technically infinite token generation. Park et al. (2024) make a similar observation and propose an adaptive algorithm wherein a beam search process discovers valid and optimal token sequences according to the grammar. They report some success in generating valid sequences, but note accuracy is not improved in all respects. Finally, Beurer-Kellner et al. (2024) propose “DOMINO,” a “minimally invasive constrained decoding” algorithm for solving the alignment problem by pre-computing allowed sequences of tokens before dynamically selecting certain paths according to probability mass. While all three approaches show promise, we note all propose non-arbitrary adjustments to GCD as is commonly presently implemented. This limits their downstream applicability and leaves open the question of whether there exists a generalizable set of token alignment strategies useful for common inference servers as are presently implemented. In contrast, this paper evaluates strategies for selecting better grammars within existing systems.
@@
Artifacts in LLM training can result in tokens that appear semantically similar to humans but whose latent representations differ to a great extent. Recent literature has drawn attention to how LLMs capture semantic similarity in the embedding space. Wen-Yi and Mimno (2023) explore trained token representations in the embedding weights of highly multilingual LLMs, observing significant variation in how models represent semantically similar tokens in the embedding space: worryingly, they find certain models fail to represent shared semantics between otherwise closely related tokens. The authors leave open the question of whether asymmetric token embeddings negatively impact downstream performance. Zhang et al. (2024) report similar behavior in monolingual LLMs, noting asymmetries in token categories including pronouns, numbers, and nouns. They likewise do not measure the downstream impact of this behavior.
@@
Researchers are similarly investigating the downstream consequences of models whose tokenizers incorporate leading whitespace (henceforth LW) into word-leading subword tokens. Tokenizers descending from SentencePiece duplicate certain tokens: once with LW to capture instances where the token begins a word in languages making use of the space, and once without (Kudo and Richardson, 2018). For example, the Llama 3.1 tokenizer can represent the word “culture” either as the token culture (index key $70905$) or _culture (index key $7829$). These are two distinct tokens despite appearing identical to humans once decoded by the tokenizer. The goal of this duplication is efficiency, in that frequently issuing a dedicated space token when tokenizing regular English text can increase (in the worst case) token counts by a count of $12n-$ 1 where $n$ is the number of words in the sequence. There are other consequences. Pimentel and Meister (2024) note tokenizers making use of LW tokens can complicate computing token surprisal given model likelihoods are attributed on a per-subword basis. Oh and Schuler (2024) similarly find models trained with LW tokens are prone to an increased likelihood of failing to learn word boundaries given their tokenizers explicitly penalize the appearance of space tokens. Neither work estimates the impact LW tokens may have on downstream tasks, nor whether LW tokens can eventually semantically converge with non-LW tokens during pre-training.
@@ 3 Methods
To mitigate grammar-model misalignment on existing GCD implementations, we compare output formats. The phenomenon of non-exchangeable token embeddings suggests that two formats that appear identical to humans could have different performance. Can we dependably elicit these preferences by constructing grammars with particular sets of tokens? If so, we can avoid extending existing GCD-enabled inference servers as is commonly recommended by the literature (Geng et al., 2023; Park et al., 2024; Beurer-Kellner et al., 2024). In this section we describe our method for systematically discovering these GCD-optimal token sets.
Description Token Set Max. Length Example
Integer
Integer, numeric $\{1,2,...,10\}$ 1 token 7
@@
Real Number
Real, numeric $\{0.1,...,1\}$ $\leq 3$ token(s) 0.75
Real, word {Zero point zero, …, One} $\leq 4$ token(s) “Zero point seventy five”
Percentage
@@
Percentage, word {Zero, …, One hundred; percent} $\leq 4$ token(s) “Seventy five percent”
Boolean
Binary, numeric {1, 0} 1 token 1
@@
Likert
Likert task, numeric $\{1,2,3,4,5\}$ 1 token 5
Likert task, word {Strongly disagree, …, Strongly agree} $\leq 3$ token(s) “Strongly agree”
Figure 2: Target token sets considered in our study, noting the value range and specifying the maximum length of a valid sequence with Llama 3.1 followed by an example.
@@
We assess the viability of certain token sets against four benchmarks. Our first two benchmarks are the Semantic Textual Similarity Benchmark (Cer et al., 2017) and the MEN Test Collection (Bruni et al., 2014). Both are semantic textual similarity (STS) tasks asking annotators to capture the semantic similarity of two sentences on an interval scale uniform between two numbers. Our next two benchmarks are Quora Question Pairs (DataCanary et al., 2017) and ToxicChat (Lin et al., 2023). Both are detection tasks asking annotators to assess whether a given phenomenon is present; their annotation target is a binary. We randomly sample $4,000$ problems from across all benchmarks with equal numbers from each. We then specify a task prompt template valid for problems sampled from all four datasets:
<string 1>{word_1}</string 1>
<string 2>{word_2}</string 2>
Rate your agreement with the following
statement: {statement}
{constraint}
The template accepts four parameters. Two (word_1 and word_2) are problem-specific, in that all benchmarks construct problems in the form of a pair-wise string comparison. statement is benchmark-specific, in that the fundamental task posed by each benchmark is summarized in a single sentence (e.g. These strings are similar.). Finally, constraint contains a short instruction informing the model which token sets are permissible. We present example task prompts in Appendix A.
@@
We contextualize our LLM results using two baseline measures. Since two of our four benchmarks target semantic similarity as a real number, we make use of BERTScore to provide a baseline BERT-based cosine similarity score for every problem in all four benchmarks (Zhang et al., 2020). This system is specifically tuned to perform semantic similarity scoring, and serves as an upper bound relative to zero-shot prompting strategies. We note the semantic similarity problems in Bruni et al. (2014) and Cer et al. (2017) together with the duplicate questions in DataCanary et al. (2017) all appear lightly correlated with string edit distance, so we also use Levenshtein distance to provide a non-LLM lower bound for all problems across all benchmarks (Levenshtein, 1966). We intend neither measure to be state-of-the-art. Rather, the BERTScore measures to represent the performance of a reasonable language model, while the edit distance of each problem is a simple method that is still expected to outperform random guessing.
@@
To validate whether differences between grammars are consistent or model-dependent, we evaluate on eight models of varying size belonging to four popular families of instruction-tuned LLMs: Llama 3.1 8b/70b, Qwen 2.5 7b/72b, Gemma 2 9b/24b and Phi 3 3b/14b (Dubey et al., 2024; Qwen et al., 2025; Team et al., 2024; Abdin et al., 2024). All models were trained in the last two years and feature pre-training sets ranging between $3.3$ and $18$ trillion tokens. We select these model families for three reasons: their high performance relative to proprietary models, their high compatibility with GCD-enabled inference servers, and their popularity with researchers for certain downstream tasks such as classification and annotation, e.g. Bamman et al. (2024).
@@
All models are run with default temperature and Top_p. Context lengths are set to $512$ tokens. We obtain models from their official repositories and quantize all models to q8_0 with scripts provided by Gerganov (2024) to maximize local inference speed. We follow prior research on quantized models suggesting that reducing weight precision to an average of no fewer than eight bits ensures performance remains comparable to unquantized models (Dettmers et al., 2022). We host the models with llama.cpp (compiled with Cuda support) and run the server on a system equipped with two Nvidia RTX A6000 GPUs. We make use of use the GBNF GCD implementation included with llama.cpp without modification. [^1] Examples of GBNF grammars can be found in the llama.cpp docs.
@@
In current GCD implementations, users specify output formats either with grammars or regular expressions. We can make arguments for several formats. A fine-grained floating point number may provide more flexibility in output, but a model could also do better with less flexibility and a coarser, more restricted output space like a binary or Likert scale.
We consider five commonly used output formats, shown in Figure 2. These categories include integers between $1$ and $5$ inclusive, real numbers in the closed interval $[0,1]$, percentages, binaries, and $5$-point Likert scales. In addition to the output domain, we add three additional format variables:
@@
We vary across datasets, token families, model families, model sizes, treatments, and our baseline measures. These combinations result in $1,280,000$ total classifications.
@@ 4 Results
We first evaluate the two baseline measures (BERTScore and Levenshtein distance) over all four benchmarks. As expected, BERTScore, a system specifically trained for semantic similarity, shows strong correlation on two semantic similarity benchmarks ($r\!=\!0.755$ and $r\!=\!0.839$). It performs less well on Quora Pairs $r\!=\!0.540$ and near random on ToxicChat $r\!=\!-0.061$. These values serve as an indication of the capabilities and limitations of systems fine-tuned for particular tasks. On the other extreme, the Levenshtein distance baseline performs poorly in all cases, with mean correlation of $r\!=\!-0.094$. This result replicates Lin et al. (2023), whose task involves classifying the presence of toxic language in textual transcripts and is thus explicitly not measuring semantic similarity. This result indicates that the tasks are indeed hard and cannot be solved by naïve algorithms.
Real numbers perform best.
Correlations (Small Models)
Model Integer Real Percent Binary Likert
Llama .$-.160$$420$$046$$199$191.191.191.191
Gemma .$181$$\mathbf{.537}$$346$$118$138.138.138.138
Phi .$436$$\mathbf{.505}$$422$$402$453.453.453.453
Qwen .$331$$\mathbf{.568}$$372$$344$466.466.466.466
$\mathbf{\bar{x}}$$197$$507$$296$$266$312.312.312.312
Correlations (Large Models)
Model Integer Real Percent Binary Likert
Llama .$151$$492$$077$$306$311.311.311.311
Gemma .$155$$476$$110$$095$$-.006$
Phi .$362$$412$$376$$335$413.413.413.413
Qwen .$429$$\mathbf{.609}$$\mathbf{.659}$$472$$\mathbf{.581}$
$\mathbf{\bar{x}}$$274$$497$$306$$302$325.325.325.325
[…]
Table 1 shows correlations between model output and human labels broken down by model and token set, but not adding any additional formatting such as newline after the beginning of the assistant turn or leading space before the classification token. Results are the mean of observations across all benchmarks. Real numbers perform best across all models, with a minimum Spearman rank correlation coefficient of $\rho\!=\!0.420$ and a maximum of $\rho\!=\!0.609$. Conversely, we find integers and binaries perform worst, particularly, $\rho\!=\!-0.160$ for integers on Llama 3.1 8b. Likert proves the least reliable, with the label range yielding a correlation of $\rho\!=\!-0.006$ with Gemma 2 27b and $\rho\!=\!0.581$ using Qwen 2.5 72b. These results indicate that there are consistent, substantial differences between output formats, and that real-valued outputs are consistently the best format across all models. We offer further exploration in section 5.
@@
Averaging over all formats, large models have higher correlation than small models, at $0.341\pm 0.28$ and $0.316\pm 0.39$ respectively. This trend is due mostly, however, to the Llama and Qwen families. Phi 3 and Gemma 2 show little improvement or even reduction in performance. For Likert scores, Gemma goes down from $\rho\!=\!0.138$ to $\rho\!=\!-0.006$ between small and large models. We investigate this phenomenon more closely in section 5. The difference between the real-valued output format and the other formats is also less pronounced for large models than for small models. The global mean values likewise disguise variations in per-token set performance. While increasing parameter count improves correlation for most formats, we find integers exhibit a mean increase of less than 3.5%. This intra-range variation is most pronounced in Gemma 2 27b, which yields lower mean correlation scores for the percent, binary, and Likert token sets.
@@
Correlations (Post-Treatment)
@@
as_large .$\mathbf{.320}$$048$$069$093.093.093.093
with_newline .$\mathbf{.250}$$\mathbf{.390}$$-.110$$\mathbf{.130}$
with_space .$\mathbf{.140}$$096$$034$$\mathbf{.130}$
as_word .$-.180$$-.180$$-.410$$-.120$
format .$-.180$$-.130$$-.028$$-.180$
[…]
In addition to the size of the model and the output domain, we also vary several additional format variables, as shown in Table 2. Using digits rather than words as_word reduces performance. We find all models exhibit strong preferences for digit tokens when completing tasks involving quantities and selection (e.g. assigning confidence or assessing similarity). Asking for numeric values as English words reduces correlation for all models: $\rho\!=\!-0.180$ for Llama, $\rho\!=\!-0.180$ for Gemma, $\rho\!=\!-0.410$ for Phi, and $\rho\!=\!-0.120$ for Qwen. Format likewise does not result in a significant change to performance when averaged over all observations.
We see substantial improvement from increasing parameter count as_large, as expected. However, including an initial new line with_newline (including leading whitespace in the token immediately prior), and with_space (adding leading whitespace in the token itself). As previously described, increasing parameter count is a dependable method for increasing downstream performance. That with_newline and with_space reliably increase performance across the Llama, Gemma, and Qwen families is a more suggestive result: both involve adding whitespace. The first treatment, with_newline, involves adding a newline following the beginning of the assistant turn in most chat templates. The second, with_space, involves the use of tokens directly encoding leading whitespace. Both increase performance, but only one is necessary for the effect as we find the presence of with_newline causes with_space to lose statistical power. Moreover, we observe neither Phi 3 model respond to LW. We investigate this further in section 5.

### v2

# Lost in Space: Finding the Right Tokens for Structured Output

@@ Abstract
General-purpose language models are trained to produce varied natural language outputs, but for some tasks, like annotation or classification, we need more specific output formats. LLM systems increasingly support structured output, which enforces formats by sampling tokens according to a grammar — but also unpredictably reduces downstream performance. Are there systematic differences between grammars that appear semantically (and often visually) similar to humans? To answer this, we test four popular model families with five varying output formats on four common NLP benchmarks. We find all models perform most accurately when guided to use formats respecting convention, such as letters for multiple choice and real numbers for numerical prediction. Performance also improves by 5%-10% when guiding models to return tokens incorporating leading whitespace, with smaller models benefiting the most. We find leading whitespace helps models avoid structural deficiencies in subword token representations. We finally present best practices for researchers using language models as zero-shot classifiers with structured output.
Lost in Space: Finding the Right Tokens for Structured Output
@@ 1 Introduction
Prompt: Rate the similarity of these words from 1 to 5: “train” and “underground”. Use only the following characters:
Tokens decoded as digits: 12345         → 3
Tokens decoded as digits with whitespace: ␣1␣2␣3␣4␣5 → 5
Llama 3.1 picks different tokens depending on whether given tokens incorporate whitespace.
@@
Figure 1: Llama 3.1 8b provides different answers to a question from the MEN Test Collection when constrained to produce digit tokens with and without a leading space. Sampled with temperature set to 0.
Language model outputs are flexible, which makes them effective for a wide range of applications, but this can also be a liability. While language models are often used to emit free text, e.g. writing a story, other important applications require their output match specific formats like class labels and domain-specific languages (Wang et al., 2023a; Geng et al., 2023). Rigid formats make it easier to parse responses, helping users avoid wasting queries on potentially invalid stochastic output.
@@
Multiple choice options are the simplest example, but users are increasingly using more flexible (and more failure-prone) output format restrictions. Reinforcement learning from human feedback (RLHF) and instruction tuning can guarantee well-formed output (Ouyang et al., 2022), but these methods are time consuming and expensive to implement, and are impractical for the variable and rapid-turnaround scenarios typical of LLM usage.
@@
An increasingly popular method for ensuring output consistency without fine-tuning is selectively sampling tokens valid according to a formal language (Shin et al., 2021; Wang et al., 2023b). Structured output, powered by grammar-constrained decoding (GCD), is regularly implemented in server software for running LLM inference (Gerganov, 2024; Microsoft, 2024; Rickard, 2024). Some servers allow users to specify grammars with regular languages (Willard and Louf, 2023; Microsoft, 2024), while others allow users to write context-free grammars in Backus–Naur form (Chomsky, 1956; Knuth, 1964; Gerganov, 2024).
But despite this popularity, enabling GCD appears to degrade downstream task accuracy (Willard and Louf, 2023; Park et al., 2024; Beurer-Kellner et al., 2024; Geng et al., 2023; Le et al., 2024). We hypothesize that the reduced performance of GCD is not due to GCD itself, but to non-optimal grammars. For example, Figure 1 shows Llama 3.1 produces different results when constrained by two different output token sets, one with leading spaces and another without.
@@
We investigate this phenomenon by comparing the performance achieved when using one of five distinct token sets across four different data labeling scenarios with four popular families of pre-trained language models. We present three key contributions. First, although all models vary considerably between formats for a given task, they often agree on a preferred format, such as real numbers for prediction tasks. Second, important distinctions may not be obvious to humans. We find that tokens incorporating leading whitespace consistently improve performance, echoing results for syntactic tokens (Pimentel and Meister, 2024; Oh and Schuler, 2024). We offer initial explanation for this difference using evidence in the embedding weights of our models of interest. We finally propose a set of strategies for selecting target tokens when specifying grammars for GCD libraries. These strategies can help achieve better performance when using structured output, particularly for users of smaller, less computationally intensive models.
@@ 2 Related Work
Grammar-model misalignment is a recognized issue with grammar-constrained decoding algorithms. Existing solutions generally involve modifying GCD algorithms to avoid misalignment by adding additional constraints. Park et al. (2024) propose an adaptive algorithm where beam search discovers valid and optimal token sequences according to the grammar. They report some success in generating valid sequences, but note accuracy is not improved in all respects. Beurer-Kellner et al. (2024) propose “DOMINO,” a “minimally invasive” algorithm, pre-computing allowed sequences of tokens before dynamically selecting certain paths according to probability mass. While these approaches show promise, both propose non-trivial adjustments to how GCD is commonly implemented in software. This limits their downstream applicability, and does not provide a generalizable token alignment strategy for common inference servers as they are presently implemented.
@@
Artifacts in LLM training can result in tokens that appear semantically similar to humans but whose latent representations differ to a great extent. Wang and Mimno (2023) explore trained token representations in the embedding weights of highly multilingual LLMs, observing significant variation in how models represent semantically similar tokens in the embedding space. Zhang et al. (2024) report similar behavior in monolingual LLMs, noting asymmetries in token categories such as pronouns and nouns. Neither study assesses the downstream impact of this behavior.
@@
Tokenizers descending from SentencePiece duplicate certain tokens: once with leading whitespace to capture instances where the token begins a word, and again without whitespace (Kudo and Richardson, 2018). For example, the Llama 3 tokenizer can represent “culture” either as the token culture (index key 709057090570905) or ␣culture (782978297829). These tokens are distinct despite decoding to strings nearly identical to humans.
Pimentel and Meister (2024) note tokenizers making use of LW tokens can complicate computing token surprisal. Oh and Schuler (2024) similarly find models trained with LW tokens are prone to an increased likelihood of failing to learn word boundaries, given their tokenizers explicitly penalize the appearance of space tokens. Neither work estimates the impact tokens may have on downstream tasks, nor whether LW tokens can converge with non-LW tokens in training.
@@ 3 Methods
If we can dependably elicit grammar preferences, we can avoid extending existing GCD-enabled inference servers as is routinely recommended by the literature (Geng et al., 2023; Park et al., 2024; Beurer-Kellner et al., 2024). In this section we describe our method for systematically discovering these GCD-optimal token sets.
Category Variant Token Set Maximum Length Example
[…]
@@
[…]
@@
[…]
Table 1: Target token sets considered in our study, noting the value range and specifying the maximum length of a valid sequence with Llama 3.1 followed by an example.
@@
We assess the viability of certain token formats against four widely-used NLP benchmarks that involve quantitative outputs. [^1] Common multiple choice benchmarks (e.g. MMLU) are subject to token influence (see Table 5), but do not require GCD as responses can be read straight from model logits. Compare with prediction tasks where responses may be composed of two to three tokens, e.g. 12% being composed of the tokens [1], [2], [%]. Our first two benchmarks are the Semantic Textual Similarity Benchmark (Cer et al., 2017) and the MEN Test Collection (Bruni et al., 2014). Both are prediction tasks asking annotators to rate the semantic similarity of two sentences on a scale between two numbers. Our next two benchmarks are binary detection tasks asking annotators to assess whether a given phenomenon is present: Quora Question Pairs (DataCanary et al., 2017) and ToxicChat (Lin et al., 2023).
We randomly sample 1,000 problems from each benchmarks for a total of 4,000 problems. We then specify a standard task prompt template valid for problems sampled from all four datasets. We present example task prompts for all four benchmarks in Appendix A.
@@
We contextualize our LLM results using two baseline measures. Given two of our four benchmarks target semantic similarity as a real number, we make use of BERTScore to provide a baseline BERT-based semantic similarity score for all problems (Zhang et al., 2020). This system was trained to perform semantic similarity scoring, conveniently serving as an upper bound to compare our zero-shot prompting strategies against. We estimate the lower bound by measuring the Levenshtein distance (Levenshtein, 1966) between both choices in all problems, with the expectation these scores will trend lower given all benchmarks appear only lightly correlated with string edit distance.
We intend neither measure to be state-of-the-art. Rather, BERTScore represents the performance of a reasonable language model, while the edit distance of each problem is a simple method expected to outperform random guessing by some margin.
@@
To assess whether grammar-dependent degradations are consistent between LLMs, we evaluate eight models of varying size. [^2] We do not include proprietary models in our assessment because existing APIs, e.g. that of OpenAI, do not enable defining context-free grammars. We select these eight from four popular families of instruction-tuned LLMs: Llama 3.1 8b/70b, Qwen 2.5 7b/72b, Gemma 2 9b/24b and Phi 3 3b/14b (Dubey et al., 2024; Qwen et al., 2025; Team et al., 2024; Abdin et al., 2024). All models were trained in the last two years and feature pre-training sets ranging between 3.3 and 18 trillion tokens. We select these model families for three reasons: their high performance relative to proprietary models, their high compatibility with GCD-enabled inference servers, and their popularity with researchers for certain downstream tasks such as classification and annotation, e.g. Bamman et al. (2024).
@@
All models are run with model-default temperature and Top_p parameters. We set context length to 512 tokens in all cases. We obtain models from their official repositories, quantizing each to q8_0 with scripts provided by Gerganov (2024) to maximize local inference speed. [^3] Prior research suggests reducing weight precision to an average of no fewer than eight bits ensures quantized performance remains comparable to unquantized models (Dettmers et al., 2022). We host the models with llama.cpp (compiled with CUDA support) and run the server on a system equipped with two Nvidia RTX A6000 GPUs. We use the GBNF GCD implementation included with llama.cpp without modification. [^4] Examples of GBNF grammars can be found here.
@@
BERTScore Levenshtein
Benchmark Corr. MSE Corr. MSE
MEN 0.768 397 0.055-0.055- 0.055 3,736
QUORA 0.609 2,574 0.323-0.323- 0.323 8,106
STSB 0.824 379 0.107-0.107- 0.107 8,076
TOXICCHAT 0.081-0.081- 0.081 4,248 0.014-0.014- 0.014 267,889
$\mathbf{\bar{x}}$ 0.530 0.125-0.125- 0.125
Table 2: Baseline correlations and MSE for each STS benchmark considered in this paper. Note semantic similarity consistently outperforms Levenshtein distance.
Current GCD implementations have users specify output formats with CFGs meaning many possible output formats exist.
We consider five commonly used output formats, shown in Table 1. These categories include integers between 1 and 5 inclusive, real numbers in the closed interval [0,1][0,1][ 0 , 1 ], percentages, binaries, and 5-point Likert scales. We further consider three additional format variables:
@@
We vary across datasets, token families, model families, model sizes, treatments, and our baseline measures. This results in 1,280,000 classifications.
@@ 4 Results
Correlations — Small Models
Model Integer Real Percent Binary Likert
Llama 0.160-0.160- 0.160 0.420 0.0460.0460.046 0.1990.1990.199 0.1910.1910.191
Gemma 0.1810.1810.181 0.537 0.3460.3460.346 0.1180.1180.118 0.1380.1380.138
Phi 0.4360.4360.436 0.505 0.4220.4220.422 0.4020.4020.402 0.4530.4530.453
Qwen 0.3310.3310.331 0.568 0.3720.3720.372 0.3440.3440.344 0.4660.4660.466
$\mathbf{\bar{x}}$ 0.1970.1970.197 0.507 0.2960.2960.296 0.2660.2660.266 0.3120.3120.312
Correlations — Large Models
Model Integer Real Percent Binary Likert
Llama 0.1510.1510.151 0.492 0.0770.0770.077 0.3060.3060.306 0.3110.3110.311
Gemma 0.1550.1550.155 0.476 0.1100.1100.110 0.0950.0950.095 0.006-0.006- 0.006
Phi 0.3620.3620.362 0.4120.4120.412 0.3760.3760.376 0.3350.3350.335 0.413
Qwen 0.4290.4290.429 0.6090.6090.609 0.659 0.4720.4720.472 0.5810.5810.581
$\mathbf{\bar{x}}$ 0.2740.2740.274 0.497 0.3060.3060.306 0.3020.3020.302 0.3250.3250.325
Table 3: The real number format has the highest Spearman correlation with human labels. Larger models show less differentiation between formats.
@@
We first evaluate the two baseline measures (BERTScore and Levenshtein distance) over all four benchmarks. We present correlations and MSE in Table 2. As expected, BERTScore, a system specifically trained for semantic similarity, shows strong correlation on our two semantic similarity benchmarks ($r\!=\!0.755$ and $r\!=\!0.839$). It performs less well on Quora Pairs $r\!=\!0.540$ and near random on ToxicChat $r\!=\!-0.061$. These values serve as an indication of the capabilities and limitations of systems fine-tuned for particular tasks. On the other extreme, the Levenshtein distance baseline performs poorly in all cases, with mean correlation of $r\!=\!-0.094$. This result replicates Lin et al. (2023), whose task involves classifying the presence of toxic language and is thus explicitly not measuring semantic similarity. This result indicates that the tasks are indeed hard and cannot be solved by naïve algorithms.
Real numbers perform best for prediction.
Table 3 shows correlations between model output and human labels broken down by model and token set. [^5] We do not provide the corresponding MSE for our non-baseline correlations due to the inherent limited range of certain token sets, e.g. binary scores and Likert numbers, being a confounding factor. Model scores were obtained without any additional formatting, such as newline after the beginning of the assistant turn or leading space before the classification token. Results are the mean of observations across all benchmarks. [^6] We emphasize here that absolute correlation values are less meaningful than the relative scores across model types. Real numbers perform best across all models, with a minimum Spearman rank correlation coefficient of $\rho\!=\!0.420$ and a maximum of $\rho\!=\!0.609$. We conversely find integers and binaries perform worst, e.g. $\rho\!=\!-0.160$ for integers on Llama 3.1 8b. Likert proves the least reliable, with the label range yielding a correlation of $\rho\!=\!-0.006$ with Gemma 2 27b and $\rho\!=\!0.581$ using Qwen 2.5 72b.
These results indicate that there are consistent, substantial differences between output formats, and that real-valued outputs are consistently the best format across all models in prediction tasks. We offer further exploration in section 5.
@@
Averaging over all formats, large models have higher correlation than small models, at 0.341±0.280.341$\pm$ 0.280.341 ± 0.28 and 0.316±0.390.316$\pm$ 0.390.316 ± 0.39 respectively. This trend is due mostly, however, to the Llama and Qwen families. Phi 3 and Gemma 2 show little improvement or even reduction in performance. For Likert scores, Gemma goes down from $\rho\!=\!0.138$ to $\rho\!=\!-0.006$ between small and large models. We investigate this phenomenon more closely in section 5. The difference between real values and the other formats is also less pronounced for large models than for small models. The global mean values likewise disguise variations in per-token set performance. While increasing parameter count improves correlation for most formats, we find integers exhibit a mean increase of less than 3.5%. This intra-set variation is most pronounced in Gemma 2 27b, which yields lower mean correlation scores for the percent, binary, and Likert formats.
@@
Treatment — Correlation Deltas ($\Delta\rho$)
@@
with_newline 0.056 0.138 0.006 0.037
with_space 0.037 0.016 0.036 0.011
as_word -0.497 -0.510 -0.451 -0.815
as_large 0.005 -0.019 0.014 0.059
$\mathbf{\bar{x}}$0.100 -0.094 -0.099 -0.177
Table 4: Treatment effect deltas. Positive values mean treatment improves correlation with ground truth.
Benchmark Stock With Newline With Space As Integer As Real As Word
ARC-C 59 77(+30%) 63   (+7%) 32   (-46%) 22   (-63%) 54   (-8%)
BoolQ 38 58(+53%) 41   (+8%) 56   (+47%) 46   (+21%) 32   (-16%)
CommonsenseQA 39 75(+92%) 62   (+59%) 29   (-26%) 20   (-49%) 59   (+51%)
HellaSwag 43 62(+44%) 37   (-14%) 28   (-35%) 24   (-44%) 43   (+0%)
MMLU 50 62(+24%) 52   (+4%) 34   (-32%) 24   (-52%) 48   (-4%)
OpenBookQA 58 79(+36%) 66   (+14%) 36   (-38%) 23   (-60%) 66   (+14%)
Winogrande 53 58(+9%) 56   (+6%) 49   (-8%) 51   (-4%) 54   (+2%)
$\mathbf{\bar{x}}$ 49 67(+37%) 54   (+10%) 38   (-22%) 30   (-39%) 51   (+4%)
Table 5: Llama 3.1 8b’s performance on seven common reasoning benchmarks after format and treatments have been applied. As predicted, leading whitespace positively impacts response accuracy. Bold values indicate best performance per benchmark. Percentages show change relative to stock baseline.
To capture what other factors matter beyond model size and token set, we vary several additional format variables described in section 3. We calculate the values in Table 4 by taking the delta between each model family’s correlation with and without the treatment, averaged across all benchmarks. A positive delta indicates the treatment improved model performance relative ground truth. This process reveals a number of key trends. First, as_word dramatically reduces performance. That is, enforcing numeric values as English words reduces correlation for all models: $\Delta\rho\!=\!-0.497$ for Llama, $\Delta\rho\!=\!-0.510$ for Gemma, $\Delta\rho\!=\!-0.451$ for Phi, and $\Delta\rho\!=\!-0.815$ for Qwen. Second, although we do find an improvement from increasing model parameter count as_large, this improvement is less significant than including either with_newline (leading whitespace in the token immediately prior), and with_space (leading whitespace in the token itself). Doing so results in an even greater performance increase across the Llama, Gemma, and Qwen families. We discuss this phenomenon further in section 5.